Scott's Winterizer Lawn Fertilizer is 22% nitrogen. Schultzs Lime with Lawn Fertilizer is 4% nitrogen. William Weaver, owner of Weaver's Nursery, wishes to mix these two fertilizers to make 400 pounds of a special 10% nitrogen mixture for midseason lawn feeding. How much of each fertilizer should he mix?
Answer by mananth(16949)

percent ----------------quantity
Scott's 22 ----------------x pounds
Schultz 4 ----------------400-x pounds
Total 10.00% ----------------400
...
22x+4(400-x)=10 *400
22x+1600-4x=4000
22x-4x=4000-1600
18x=2400
/18
x=133.33 pounds 22.00%
Balance 266.67 pounds 4.00%